OpenStudy (anonymous):

graph f(x)=4x^2-8x+1 starting with the graph y=x^2

hero (hero):

You can do that, but if I were you, I would re-write 4x^2 - 8x + 1 in vertex form first.

OpenStudy (anonymous):

whats the vertex form?

hero (hero):

f(x) = a(x-h)^2 + k, where (h,k) is the vertex. In this case: f(x) = 4(x-1)^2 - 3

hero (hero):

So now you can properly graph starting with y = x^2

OpenStudy (anonymous):

where did the 3 come from?

OpenStudy (anonymous):

@Hero

hero (hero):

find the vertex of f(x) = 4x^2 - 8x + 1, then maybe you'll understand

OpenStudy (anonymous):

i got it :) thank you @Hero

hero (hero):

Are you done?

hero (hero):

Did you start with y = x^2 and then get to f(x) = 4x^2 - 8x + 1